OpenStudy (anonymous):

Trig: How does this equal one? \( tan^2 \theta - sec^2 \theta = 1 \)

OpenStudy (anonymous):

I have \( tan^2 \theta - sec^2 \theta = 1 \) \( \frac{sin^2 \theta}{cos^2 \theta} - \frac{1}{cos^2 \theta} \) I am stuck here

OpenStudy (anonymous):

I think I got it

OpenStudy (astrophysics):

It shouldn't

OpenStudy (anonymous):

Let me know if this is correct \( tan^2 \theta - sec^2 \theta = 1 \) \( \frac{sin^2 \theta}{cos^2 \theta} - \frac{1}{cos^2 \theta} \) \( \frac{1+cos^2 \theta}{cos^2 \theta} - \frac{1}{cos^2 \theta} \) \( \frac{cos^2 \theta}{cos^2 \theta} = 1 \)

OpenStudy (freckles):

do you mean -1? recall \[\sin^2(\theta)+\cos^2(\theta)=1 \] divide both sides by cos^2(theta) (this equation comes )